GlueX TOF Meeting, November 10, 2011
Thursday, November 10, 2011
10:00 am EDT
JLab: CEBAF Center, Room F326
- Minutes from the last meeting
- Mechanical design: Tim, Chuck
- Prototype status: Paul, Sasha
- Contract Status: Mark
Talks can be deposited in the directory
/group/halld/www/halldweb/html/talks/2011-4Q on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2011-4Q/ .
Present: FSU: Sascha Ostrovidov JLab: Eugene Chudakov, Chuck Hutton, Mark Ito (chair), Beni Zihlmann
Chuck presented the current status of drawings on how to support the counters, in particular how to secure the light guides, phototubes, and shielding sleeves. Much more realism has been added, among other features, a set of collars to attach the steel shielding to the support structure.
Mark reported that we have drafts of all documents needed to start the procurement process: statement of work, sole source justification, specifications for both PMT's and scintillator, and drawings. There are still some edits to do, however, and nothing has been signed yet.
Sasha reported work in three areas.
Non-mixed set of splitters
Measurements were done where the splitters were the same on all six channels of the prototype. He tried two configurations: a set of 50/50 splitters and a set of 80/20 splitters. He measured the following time resoltions for mean time in a single layer:
Sascha tried pulsing the electronics using a pulser in place of the phototube. No reflections were observed. He still sees reflections from phototube pulses with the standard 50 foot cable between the tube and the splitter, but if another cable is added to give a total of 100 feet, the reflections are not present. He notes that in the latter case, the original and reflection are clearly separated in time. Not clear what is going on here.
Sacha broke out the 8-stage tubes that FSU bought some time ago. With a test scintillator and an americium source he sees response with about a factor of 4 or 5 times less gain than with the 10-stage tubes, as expected, consistent with a gain of 5×105 as advertised on the data sheet.
When looking at cosmics with the test scintillator, he does not see clear separation of pedestal and signal. He will try an amplifier on the output to see if that helps. Some details, for the record:
- pedestal at 300 counts
- signal at 400-500 counts
- 150 ns gate
- 0.25 pC/count